2016 part c

--

Please calculate the following ratios

Price per share = 83.95

Return on Asset (ROA)

Net Profit Margin

----------

Price to Earnings ratio

Book-to-market Ratio



CONSOLIDATED STATEMENT OF STOCKHOLDERS' EQUITY-USD (S) shares in Millions, $ in Millions Common Stock In Common Stock In Accumulated Other Total Cameron Common Stock Common Stock Balance at Dec. 31, 2013 Balance, shares at Dec. 31, 2013 Treasury shares, Balance at Dec. 31 S 12,192 1,434 $ (8,135) S 37,966 S 166 $ 5,506 5,438 Changes in unrealized gain on Changes in fair value of cash fiow Shares sold to optionees, less shares (26) S 556 Shares sold to optionees, less shares (79) Vesting of restricted stock, shares Shares issued under employee stock $ 295 262 Shares issued under employee stock program (4,678) Stock repurchase program, shares (2,071) Shares issued for acquisition, shares Balance at Dec. 31, 2014 Balance, shares at Dec. 31, 2014 Treasury shares, Balance at Dec. 31 S 12,495 1,434 11,772) 41,333 (4.206) 199 $ 2,135 2,072 Changes in unrealized gain on Changes in fair value of cash fiow Pension and other postretirement benefit Shares sold to optionees, less shares 123 (38) S 190
Answer #1

1. The return on assets ratio formula is calculated by dividing net income by average total assets. This ratio can also be represented as a product of the profit margin and the total asset turnover. Either formula can be used to calculate the return on total assets.

Net Income given = (1,627)

Average Assets = Information Not Available, Balance Sheet required

Return on Assets = Net Income / Average Assets = NA as the Profit is Negative.

2. The Net Profit margin is calculated from the income statement, subtract the cost of goods sold,operating expenses, other expenses, interest (on debt), and taxes from revenue.

Profit = (1,627)

Revenue = 27,810

Net Profit Margin = Profit / Revenue = NA as the profit is negative.

3. The basic Price to Earning (P/E) formula takes current stock price and EPS to find the current P/E. EPS is found by taking earnings from the last twelve months divided by the weighted average shares outstanding.

EPS = (1.24)

Price = 83.95

Price to Earning = Price/ Earnings = NA as the EPS is negative

4. The book-to-market ratio is used to find a company's value by comparing its book value to its market value. A company's book value is calculated by looking at the company's historical cost, or accounting value.

Book Value = (41,529+35,905)/2 = 38,717

Book Value Per Share = Book Value / No. of Shares Outstanding = 38,717 / 1,391 = 27.833

Market Value = 83.95

Book to Market Ratio = Market Value / Book Value = 83.95 / 27.833 = 3.01